A circular disk of radius 2m is at rest and then begins to rotate with a constant angular acceleration of 200 rad / s ^ 2. Find the acceleration of a point on the circumference of the disk at the moment when it rotates 1 revolution per second.


1
Expert's answer
2021-06-10T10:22:59-0400
"a_n=\omega ^2r\\a_n=(2\pi(1))^2(2)=78.96\frac{m}{s^2}\\a_t=\alpha r=(200)(2)=400\frac{m}{s^2}\\a=\sqrt{400^2+78.96^2}=407.7\frac{m}{s^2}"